OpenStudy (anonymous):

PLEASE HELP SOLVE THIS STUDYING FOR A TEST TOMORROW AND DONT UNDERSTAND. The arch of the Sidney harbor bridge is approximately 500 meters long and 85 meters high. what quadratic function models the curve of the arch? assume the arch starts at (0,0)

OpenStudy (anonymous):

The answer is... I Dont Know

OpenStudy (anonymous):

Any curve can be made to pass through three points. So first identify the conditions you know. 1) At x = 0, y = 0 (Arch starts at 0, 0) 2) At x = 250, y = 85 (Arch is 85 meters tall) 3) At x = 500, y = 0 (The arch is 500 meters in length) From this, you know the quadratic will be curving down, so the y term will be negative. Now put in the general form of your quadratic equation... \[(y-a)^2=(x-b)^2+c\] Since you have three unknowns and three conditions, you can solve for the constants one at a time. (There is also a shortcut somewhere, but I don't remember what it would be.)

OpenStudy (anonymous):

@gypsy1274 can you please draw the picture again it was really helpful

OpenStudy (anonymous):